# AgentThink: A Unified Framework for Tool-Augmented Chain-of-Thought Reasoning in Vision-Language Models for Autonomous Driving

### v2

@@ 2.1 Language Models in Autonomous Driving
Recent works have augmented LLMs with prompting strategies, memory-based context construction, or vision inputs [12]. For instance, DriveVLM [35] introduces a CoT approach with modules for scene description, analysis, and hierarchical planning, while DriveLM [33] focuses on graph-structured visual question answering. EMMA [13] demonstrates how multimodal models can directly map raw camera inputs to driving outputs, including trajectories and perception objects. Despite these advancements, both LLM-centric and VLM-based methods often treat reasoning as a static input-output mapping, with limited ability to detect uncertainty, perform intermediate verification, or incorporate physical constraints [15]. Challenges such as hallucinations, over-reliance on rigid templates, and a lack of domain-specific reward feedback persist. To address these limitations, our work introduces a Tool-Augmented, RL-based reasoning framework that enables dynamic and verifiable decision-making for autonomous driving.

### v3

@@ 1 Introduction
"A gentleman is not different by nature, he is good at making use of things around him." - Hsun Tzu
@@ 2.1 Language Models in Autonomous Driving
Recent works have augmented LLMs with prompting strategies, memory-based context construction, or vision inputs [12]. For instance, DriveVLM [35, 28] introduces a CoT approach and dual system with modules for scene description, analysis, and hierarchical planning, while DriveLM [33] focuses on graph-structured visual question answering. EMMA [13] demonstrates how multimodal models can directly map raw camera inputs to driving outputs, including trajectories and perception objects. Despite these advancements, both LLM-centric and VLM-based methods often treat reasoning as a static input-output mapping, with limited ability to detect uncertainty, perform intermediate verification, or incorporate physical constraints [15]. Challenges such as hallucinations, over-reliance on rigid templates, and a lack of domain-specific reward feedback persist. To address these limitations, our work introduces a Tool-Augmented, RL-based reasoning framework that enables dynamic and verifiable decision-making for autonomous driving.